What is the absorbance for a sample with 79 %T? (Write answer to three decimal places, for example answer should be written as 0.366)
Answer #1

Given

percent transmittance (%T) = 79%

We know the formula

Absorbance = 2 - log(%T)

hence

Absorbance = 2 - log(79)

= 2 - 1.897

= 0.103 absorbance units

So our answer is  0.103 absorbance units